OpenStudy (xenesis98):

What are the values of x and y in this linear system? 2x + 3y = 8 x + 2y = 12

OpenStudy (sohailiftikhar):

just want answer or explanation ?

OpenStudy (sohailiftikhar):

to solve it first eliminate one variable from any equation and find the value of other variable

OpenStudy (sohailiftikhar):

so what variable do you want eliminate first ?

OpenStudy (dinamix):

2x+3y=8 2x+4y=24 2x+3y-(2x+4y)=-16 -y=-16 y=16 u can find x now @xenesis98

OpenStudy (sohailiftikhar):

ok i'm eleminating y so 2x + 3y = 8 x + 2y = 12 this the equation now multiply equation (2) with 2 you will get 2x + 3y = 8 2x + 4y = 24 now subtract eq (1) from eq(2) you will get y=16 now put this value of y=16 in any equation to find the value of x

OpenStudy (sohailiftikhar):

didn't understand ?